OpenStudy (anonymous):

Simplify √405 - 3√45?

OpenStudy (acxbox22):

\[\sqrt{405}=\sqrt{81\times5}=9\sqrt{5}\] \[-3\sqrt{45}=-3\sqrt{9\times5}=-9\sqrt{5}\] \[9\sqrt{5}-9\sqrt{5}=???\]

OpenStudy (anonymous):

Thank you! @acxbox22

OpenStudy (acxbox22):

anytime